If a weather station uses electromagnetic waves with a frequency of 2.85 x 10^9 Hz, what is the length of the wave
1
Expert's answer
2020-04-30T10:37:07-0400

As "c=\nu\lambda"

So, "\lambda=\frac{c}{\nu}=\frac{3\times10^8}{2.85\times10^9}=0.1052m"